A house that is assessed for 12,00 pays $960 in realty taxes. what should be the realty tax on a house that is assessed for $16,500?
Answer by checkley71(8403)

I'LL ASSUME YOU MEAN 12,000 FOR THE FIRST ASSESSMENT.
12000/960=16500/X
12000X=960*16500
12000X=15,840,000
X=15,840,000/12000
X=$1,320 IS THE TAX ON THE $16,500 HOUSE.